# Compelling ReLU Network Initialization and Training to Leverage Exponential Scaling with Depth

### v2

@@ 3.1 Differentiability of Model Output
$F^{\prime}(x)$ is continuous only if the scaling coefficients are selected based on $a_{i}$ according to:
@@
Proof.
Rewriting Equation (4) (which is equal to 0) for layers $i$ and $i+1$ in the following way:
$s_{i}(1-a_{i+1})=s_{i+1}+\frac{1}{a_{i+2}}\left(s_{i+2}+\sum_{n=i+3}^{\infty}s_{n}\prod_{k=i+3}^{n}\frac{1}{a_{k}}\right)$
$s_{i+1}(1-a_{i+2})=s_{i+2}+\sum_{n=i+3}^{\infty}s_{n}\prod_{k=i+3}^{n}\frac{1}{a_{k}}$
allows for a substitution to eliminate the infinite sum
$s_{i}(1-a_{i+1})=s_{i+1}+\frac{1-a_{i+2}}{a_{i+2}}s_{i+1}$
Collecting all the terms gives
$s_{i+1}=\frac{s_{i}(1-a_{i+1})}{1+\frac{1-a_{i+2}}{a_{i+2}}}$
which simplifies to the desired result. ∎
@@ 4 Experiments
In this section, we implement our method to learn several convex one-dimensional curves. We present several important comparisons. To demonstrate how our networks make better use of depth, we benchmark against PyTorch’s default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both. Additionally, we fix a set of initialization points such that the output is differentiable in the limit, and so that the output produces linear segments exponentially with depth. From these identical initialization points, we compare the effect of pretraining versus gradient descent alone. Lastly, we experiment with training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability. As in related papers (Perekrestenko et al., 2018; Daubechies et al., 2022), we restrict our attention to one-dimensional examples, which are sufficient to demonstrate our proposed theory and methodology.
@@ 4.1 Experimental Setup
All models are trained using Adam (Kingma & Ba, 2017) as the optimizer for 1000 epochs to ensure convergence. The data are 500 evenly spaced points on the interval $[0,1]$ for each of the curves. Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output. The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type. The networks unique to this paper share a common set of starting locations, so that the effects of each training regimen are directly comparable. The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave. The curves are chosen to capture a variety of convex one dimensional functions. To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$. For the other approximations, the waves are subtracted. This requires the first scaling factor to be $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$. These curves are simple, and the data is noiseless and synthetic, so we do not use a train-test split for these experiments. We instead focus on learning an interpolation of the data with a low MSE to show that effective representations can be found for these functions.
@@ 4.2 Numerical Results
Table 1: Minimum and mean ($n=30$) MSE error approximating $y=x^{3}$ and $x^{11}$.
Network Min $x^{3}$ Min $x^{11}$ Mean $x^{3}$ Mean $x^{11}$
@@
No Pretraining $7.63\times 10^{-7}$ $1.86\times 10^{-5}$ $3.89\times 10^{-5}$ $3.56\times 10^{-4}$
@@
Table 2: Minimum and mean ($n=30$) MSE error approximating $y=\sin(x)$ and $y=\tanh(3x)$.
Network Min $\sin(x)$ Min $\tanh(3x)$ Mean $\sin(x)$ Mean $\tanh(3x)$
@@
No Pretraining $1.96\times 10^{-7}$ $1.07\times 10^{-6}$ $1.93\times 10^{-5}$ $8.38\times 10^{-5}$
@@
Our main results are shown in Tables 1 and 2, wherein we observe several important trends. First, the worst performing networks are those that rely on purely randomized initializations. Even the networks that forgo pretraining benefit from initializing with many activation regions. When pretraining constraints are used, they are able to steer gradient descent to the best solutions, resulting in reductions in minimum error of three orders of magnitude over default networks.
Pretraining with differentiability enforced also closes the gap between the minimum and mean errors compared to other setups. This indicates that these loss landscapes are indeed the most reliable to traverse. When the scaling factors are trained separately, it can cause networks to exhibit behavior akin to overfitting. Even in our noiseless setting, we find that it is still possible to settle into local minima where the function interpolates only a small subset of the data points and fails to accurately represent the target curve in between them. Enforcing differentiability during pretraining can impart a bias towards smoother solutions during gradient descent and eliminates these occurrences in our experiments.
@@
Figure 5: Approximation produced with standard methodologies. The layers are shown before ReLU is applied. The neuron colors here are arbitrary, they do not correspond to Figure 2
@@ 4.3 Gradient Descent does not Directly Optimize Efficiency
Figure 5 shows the interior of a default network. The layers here are shown before applying ReLU. The default networks fail to make efficient use of ReLU to produce linear regions, even falling short of 1 bend per neuron, which can easily be attained by forming a linear spline (1 hidden layer) that interpolates some of the data points. Rather than an exponential efficiency boost, depth is actually hindering these networks. Examining the figure, the first two layers are wasted. No neuron’s activation pattern crosses $y=0$, so ReLU is never used. Layer 3 could be formed directly from the input signal. Deeper in the network, more neurons remain either strictly positive or negative. Those that intersect $y=0$ are monotonic, only able to introduce one bend at a time. The core issue is that while more bends leads to better accuracy, networks that have few bends are not locally connected in parameter space to those that have many. This is problematic since gradients can only carry information about the effects of infinitesimal parameter modifications. If a bend exists, gradient descent can reposition it. But for a neuron that always outputs a strictly positive value (such as the red in layer 2), bends cannot be introduced by infinitesimal weight or bias adjustments. Therefore, bend-related information will be absent from its gradients. Gradient descent will only introduce a network to bend by happenstance; indirectly related local factors must guide a neuron to begin outputting negative values. Rarely, these local incentives are totally absent, and the network outputs a bend-free line of best fit.
4.4 Inside the Compositional Networks
In Figure 6 we compare the top performing models with and without reparameterized pretraining. We observe that without the guidance of the pretraining, gradient descent usually loses the triangle generating structure. It typically happens around layer 4 or 5, devolving into noisy patterns and resulting in higher errors whereas pretraining maintains structure at greater depths. This behavior of gradient descent we observe is problematic since theoretical works often rely on specific constructions within networks to prove their results. Gradient descent greedily abandons any such structure in favor of models that can be worse in the long term. A theoretical result that shows a certain representation exists in the set of neural networks will thus have a hard time actually learning it without a subsequent plan to control training.
@@ 5 Limitations and Extensions
Although our reparemeterization can help maintain the triangular structure out to greater depths than unassisted gradient, our networks still lose structure eventually. This suggests that our reparameterization is limited in the functions it can represent. It may be enabling a rapid error reduction initially in the first few layers, but be unable to approach most curves to arbitrary precision. It is in this remaining error gap that the second stage gradient descent becomes lost. In the appendix, we detail two important results concerning the theoretical expressive power of our reparameterization that support this.
First, the differentiability constraints on the scaling parameters during pretraining necessitate that the network outputs a convex function. It is important for the pretraining to be able to produce a reasonably close initial solution; therefore, the reparameterization we use is only suitable for approximating convex functions. The second limitation we encounter is that pursuing further smoothness (at least in this 4-neuron configuration) is not possible. We show that approximating $x^{2}$ is the only way to have a twice continuously differentiable output in the infinite depth limit.
Despite these limitations, these networks could still be considered as replacements of the $x^{2}$ gate discussed in Section 2.1. Likewise, they could be used as a base component for assembly into larger networks. These composed networks could learn non-convex functions in higher dimensions, and they would at minimum be guaranteed to represent interpolating polynomials. The drawback is that they would likely become prohibitively large. Rather than approximating another approximation scheme out of ReLU networks, it may be better to focus on finding representations that more naturally arise from ReLU networks themselves when extending this work. Two properties specifically stand out as being somewhat fundamental to deep ReLU approximations: each linear region produced is a full copy of the original input interval $[0,1]$, and the linear regions boundaries become dense as depth tends towards infinity. Density ensures that no two inputs are linearly related in an infinite network, and the repeated reproduction of the input space allows each layer to act on each previously existing region independently and simultaneously. These properties both arise in our method, and are likely important considerations when seeking theoretical extensions. Lastly, it might be important to consider non-differentiable model outputs as a way to gain additional flexibility, but additional protections would need to be put in place against overfitting.
@@ 6 Concluding Remarks
This paper focused on exploiting the potential computational complexity advantages neural networks offer for the problem of training efficient nonlinear function representations; in particular, compelling ReLU networks to approximate functions with exponential accuracy as network depth is linearly increased. Our results showed improvements of one to several orders of magnitude in using our reparameterization and pretraining strategy to train ReLU networks to learn various nonlinear, convex, one-dimensional functions. Though further research is required to extend this technique to general multi-dimensional nonlinear functions, the present work’s principal significance lies in demonstrating the possibility of converting theoretical ReLU-based constructions into novel training procedures. This finding is particularly powerful since random initialization and gradient descent are not likely to produce an efficient solution on their own, even if it can be proven to exist in the set of sufficiently sized ReLU networks. We are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU networks and even more general types of neural networks.
Potential Broader Impact
This paper presents work whose goal is to enable more efficient neural networks. While the present work is largely theoretical, future advances in this line of research could enable the use of much smaller networks in many practical applications, which could substantially mitigate the rapidly growing issue of energy usage in large learning systems.
Appendix A Appendix
A.1 Sufficiency for differentiability
We can show that in addition to being necessary for continuity of the derivative, our choice of scaling is sufficient when $a_{i}$ are bounded away from 0 or 1.
Theorem A.1.
$F^{\prime}(x)$ is continuous if we can find $c>0$ such that $c\leq a_{i}\leq 1-c$ for all $i$.
Proof.
We begin by considering Equation (4) for layer $i$.
$s_{i}=\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right)$
Recall that this equation is telling us about the size of the discontinuities in the derivative as triangle waves are added. Each time a triangle wave is added, it can be thought of as subtracting the terms on the right. We will prove our result by substituting Equation (5) into this formula, and then verifying that the resulting equation is valid. First we would like to rewrite each $s_{n}$ in terms of $s_{i}$. Equation (5) gives a recurrence relation. Converting it to an explicit representation we have:
$s_{n}=s_{i}\left(\prod_{j=i+1}^{n}1-a_{j}\right)\left(\prod_{k=i+2}^{n+1}a_{k}\right)$ (6)
When we substitute this into Equation (4), three things happen: each term is divisible by $s_{i}$ so $s_{i}$ cancels out, every factor in the product except the last cancels, and $1-a_{i+1}$ cancels. This leaves
$1=a_{i+2}+(1-a_{i+2})a_{i+3}+(1-a_{i+2})(1-a_{i+3})a_{i+4}+...=\sum_{n=i+2}^{\infty}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})$ (7)
We will now argue that each term of the sum on the right accounts for a fraction (equal to $a_{i}$) of the remaining error. Inductively we can show:
$1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m})=\prod_{m=i+2}^{j}(1-a_{m})$ (8)
Intuitively, this means as long as the first term appearing on the right is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side. As a base case, we have $(1-a_{i+2})=(1-a_{i+2})$. Assuming the above equation holds for all previous values of $j$
$1-\sum_{n=i+2}^{j+1}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))=1-\sum_{n=i+2}^{j}a_{n}\prod_{m=i+2}^{n-1}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=$
using the inductive hypothesis to make the substitution
$\prod_{m=i+2}^{j}(1-a_{m}))-a_{j+1}\prod_{m=i+2}^{j}(1-a_{m}))=\prod_{m=i+2}^{j+1}(1-a_{m}))$
Since all $c<a_{i}<1-c$, the size of the discontinuity at the points $P_{i}$ is upper bounded and lower bounded by exponentially decaying series. Since both series approach zero, so does the series here. ∎
This lemma shows that the derivative of the finite approximation excluding $W_{i}$ is the same as that of the infinite sum.
Lemma A.2.
For all $x\in P_{i}$:
$F^{\prime}(x)=f^{\prime}_{i}(x)=\sum_{j=0}^{i-1}s_{j}W^{\prime}_{j}(x)$ (9)
Proof.
From the previous lemma we know $E^{\prime}(x)=0$. $F(x)=\sum_{j=0}^{i-1}s_{j}W_{j}(x)+E(x)$. The sum of the first $i-1$ terms are differentiable at the points $P_{i}$ since they lie between the discontinuities at $B_{i-1}$. ∎
A.2 Second Derivatives
Here we show that any function represented by one of these networks that is not $y=x^{2}$ does not have a continuous second derivative. To show this we will sample a discrete series of $\Delta y/\Delta x$ values from $F^{\prime}(x)$ and show that the left and right hand limits are not equal (unless $a_{i}=0.5$), which will imply that the continuous version of the limit for the second derivative does not exist. First we will produce the series of $\Delta x$. Let $x$ be the location of a peak at layer $i$, and let $l_{n}$ and $r_{n}$ be its immediate neighbors in $B_{i+n}$.
Lemma A.3.
If $c<a_{i}<1-c$ for all $i$, we have $\lim_{n\rightarrow\infty}r_{n}=\lim_{n\rightarrow\infty}r_{n}=x$. Furthermore, $r_{n},l_{n}\neq x_{n}$ for any finite $i$.
Proof.
Let $K$ be a constant of proportionality for the slopes on the left and right of $x$ in layer $i$. These slopes are proportional to $1/a_{i}$ and $1/(1-a_{i})$ depending on which way the triangle encompassing $x$ is oriented ($W^{\prime}_{i-1}$ could have been positive or negative at its location), We will denote these as $1/L$ and $1/R$ and reason about them later. $x_{n}$ is a peak location of $W_{i}$, so on the left side slope is negative and the right is positive. Solving for the location of $T_{i+1}(W_{i}(x))=1$ on each side will give $l_{1}=x-(1-a_{i+1})L/K$ and $r_{1}=x+(1-a_{i+1})R/K$.
On each subsequent iteration $i+n$, $x$ is a valley point and the $\Delta x$ intervals get multiplied by $a_{i+n}$. $x$ is a valley point so the left slope is positive and the right is negative, and $l_{n}$,$r_{n}$ are peak points. The slope magnitudes are given by $\frac{1}{x-l_{n}}$ and $\frac{1}{r_{n}-x}$ since $C_{i+n}$ oscillates from 0 to 1 over these spans. Solving for the new peaks again will give $l_{n+1}=x-a_{i+1}(x-l_{n})$ and $r_{n+1}=x+a_{i+1}(r_{n}-x)$. The resulting non-recursive formulas are:
$x-l_{n}=\frac{L}{K}(1-a_{i+1})\prod_{m=2}^{n}a_{i+m}\text{ and }r_{n}-x=\frac{R}{K}(1-a_{i+1})\prod_{m=2}^{n}a_{i+m}$ (10)
The right hand sides will never be equal to zero with a finite number of terms since $a$ parameters are bounded away from 00 and $1$ by $c$. ∎
Next we derive the values of $\Delta y$ to complete the proof.
Theorem A.4.
A compositional network with a second differentiable output necessarily outputs $y=x^{2}$.
Proof.
The points $l_{n}$ and $r_{n}$ are all peak locations, (9) gives their derivative values as $f^{\prime}_{i+n}(r_{n})$. Earlier we reasoned about the sizes of the discontinuities in $F^{\prime}(x)$ at $x$, since $l_{n}$ and $r_{n}$ always lie on the linear intervals surrounding $x$ as $n\rightarrow\infty$, we can get the value of $f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})$ using (8) with the initial discontinuity size set to $s_{i}(K/R)$ rather than 1. Focusing on the right hand side we get:
$f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})=s_{i}(K/R)\prod_{m=2}^{n}(1-a_{i+m})$
taking $\Delta y/\Delta x$ gives a series:
$\frac{K^{2}s_{i}}{R^{2}(1-a_{i+1})}\prod_{m=i+2}^{n}\frac{1-a_{m}}{a_{m}}$
The issue which arises is that the derivation on the left is identical, except for a replacement of $R$ by $L$. The only way for these formulas to agree then is for $R=L$ which implies $a_{i}=1-a_{i}=0.5$. ∎
A.3 Error Decay
Lastly, we show that the error of these approximations decays exponentially.
Lemma A.5.
The ratio $s_{i+2}/s_{i}$ is at most $0.25$.
Proof.
by applying formula (5) twice, we have
$s_{i+2}=s_{i}(1-a_{i+1})(1-a_{i}+2)a_{i+2}a_{i+3}$
To maximize $s_{i+2}$ we choose $a_{i+1}=0$ and $a_{i+3}=1$. The quantity $a_{i+2}-a_{i+2}^{2}$ is a parabola with a maximum of $0.25$ at $a_{i+2}=0.5$. ∎
Lemma A.6.
The function $F(x)$ is convex.
Proof.
To establish this we will introduce the list $Y^{\prime}_{i}=[F^{\prime}(V_{i}[0]),f^{\prime}_{i}(V_{i}[n]),F^{\prime}(V_{i}[2^{i}]]$, which tracks the values of the derivative at the $i$th set of valley points. All but the first and last points will have been peaks at some point in their history, so Equation (9) gives the value of those derivatives as $f^{\prime}_{i}$.
We establish an inductive invariant that the y-values in the list $Y_{i}$ remain sorted in descending order, and that $Y^{\prime}_{i}[n]\geq f^{\prime}_{i}(x)\geq Y^{\prime}_{i}[n+1]$ for $V_{i}[n]<x<V_{i}[n+1]$.
Before any of $W_{i}$ are added, $f_{0}$ is a line with derivative 0, $V_{0}$ is its two endpoints. $Y^{\prime}_{0}$ is positive for the left endpoint (negative for right) since on the far edges $F^{\prime}$ is a sum of a series of positive (or negative) slopes, Therefore the points in $Y^{\prime}$ are in descending sorted order. The second part of the invariant is true since 0 is in between those values.
Consider an arbitrary interval $(V_{i}[n],V_{i}[n+1])$ of $f_{i}$, this entire interval is between two valley points, so $f^{\prime}_{i}$ (which hasn’t added $W_{i}$ yet) is some constant value in between $Y^{\prime}_{i}[n]$ and $Y^{\prime}_{i}[n+1]$. The point $x\in P_{i}\cap(V_{i}[n],V_{i}[n+1])$ will have $F^{\prime}(x)=f^{\prime}_{i}(x)$, and it will become a member of $V_{i+1}$. This means we will have $Y_{i+1}[2n]>Y_{i+1}[2n+1]>Y_{i+1}[2n+2]$, maintaining sorted order.
Adding $s_{i}W_{i}$ takes $f_{i}$ to $f_{i}+1$ splitting each constant valued interval in two about the points $P_{i}$, increasing the left side, and decreasing the right side. Recalling from the derivation of (4) all terms but the first in the sum have the same sign, so the values in $Y^{\prime}_{i}$ are approached monotonically. Therefore on the left interval we have $Y^{\prime}_{i}[n]>f_{i+1}>f_{i}$ and on the right we have $f_{i}>f_{i+1}>Y^{\prime}_{i}[n+1]$. And so $f_{i+1}$ remains monotone decreasing. ∎
Theorem A.7.
The approximation error $E_{i}(x)$ decays exponentially.
Proof.
To get this result, we apply the previous two lemmas. Since $F$ is convex, it lies above any line segment connecting any two points on the curve. $W_{i}(x)=1$ for all $x\in P_{i}$, but $W_{i}(x)=0$ for points in $V_{i}$. Since the bend points are only of nonzero value once, $f_{i}(x)=F(x)$ for all points in $V_{i}$. $f_{i}$ is made of line segments and equals $F$ repeatedly, $E_{i}$ will be a series of positively valued curve segments. The derivative will still be decreasing on each of these intervals since it was just shifted by a constant, and each of these intervals will be convex itself.
Since $E^{\prime}(x)=0$ for $x\in P_{i}$, these points will be the locations of maximum error. Since they only have nonzero value in $W_{i}$, and $E$ is convex, $s_{i}W_{i}(x)=max(E)=s_{i}$. Since $s_{i}$ decay exponentially, we have our result. ∎

### v3

@@ 3.1 Differentiable Model Output
$F^{\prime}(x)$ is defined on $[0,1]$ only if the scaling coefficients $s_{i}$ are selected based on the triangle peaks $a_{i}$ according to:
@@ 4 Experiments
The goal of deep learning is to train a network to approximate a (generally unknown) nonlinear function. Accordingly, in this section, we implement our method to learn several convex one-dimensional curves. These nonlinear functions are known—so that we can measure error—and we remark on more difficult functions in Sections 5 and 6. The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data. To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh and Sreejith (2021), and produce errors that are orders of magnitude lower than both. For our experimental networks, we use a common set of initialization points where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation (4)). We compare the effect of reparameterized pretraining against skipping immediately to standard gradient descent. We also compare training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability. The intent of both comparisons is see if pretraining and differentiability constraints facilitate smoother navigation of the loss landscape and avoidance of local minima. Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
@@ 4.1 Experimental Setup
All models are trained using Adam (Kingma and Ba, 2017) as the optimizer with a learning rate of 0.0001 for 1,000 epochs to ensure convergence. Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output. The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type. The networks unique to this paper share a common set of starting locations, so that the effects of each training regimen are directly comparable. As in related papers (Perekrestenko et al., 2018; Daubechies et al., 2022), we restrict our attention to one-dimensional examples, which are sufficient to demonstrate our proposed theory and methodology. The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave. The curves are chosen to capture a variety of convex one-dimensional functions. To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$. For the other approximations, the waves are subtracted. This requires the first scaling factor to be $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$. The first set of data is 500 evenly spaced points on the interval $[0,1]$ for each of the curves. This is chosen to be very dense deliberately, to try to evoke the most accurate representations from each network. We determine from these tests that pretraining with differentiability enforced produces the best results, so we compare it to standard networks in our second set of experiments. We use a second set of data consisting of only 10 points, with a test set of 10 points spaced in between so as to be as far away from learned data as possible. The goal of this set of experiments is to compare the predictive capacities of the networks on unseen data.
@@ 4.2 Numerical Results
Table 1: Minimum and mean (30 samples) MSE error approximating $y=x^{3}$ and $x^{11}$.
Training Type Min $x^{3}$ Min $x^{11}$ Mean $x^{3}$ Mean $x^{11}$
@@
Pretraining Skipped $7.63\times 10^{-7}$ $1.86\times 10^{-5}$ $3.89\times 10^{-5}$ $3.56\times 10^{-4}$
@@
Table 2: Minimum and mean (30 samples) MSE error approximating $y=\sin(x)$ and $y=\tanh(3x)$.
Training Type Min $\sin(x)$ Min $\tanh(3x)$ Mean $\sin(x)$ Mean $\tanh(3x)$
@@
Pretraining Skipped $1.96\times 10^{-7}$ $1.07\times 10^{-6}$ $1.93\times 10^{-5}$ $8.38\times 10^{-5}$
@@
Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends. First, the worst performing networks are those that rely on purely randomized initializations. Even the networks that forgo pretraining benefit from initializing with many activation regions. When pretraining constraints are used, they are able to steer gradient descent to the best solutions, resulting in reductions in minimum error of three orders of magnitude over default networks.
Pretraining with differentiability enforced also closes the gap between the minimum and mean errors compared to other setups. This indicates that these loss landscapes are indeed the most reliable to traverse. Even in our noiseless setting, we find that it is still possible to settle into local minima where the function fits only a small subset of the data points and fails to accurately represent the target curve in between them. Enforcing differentiability during pretraining can impart a bias towards smoother solutions during gradient descent and eliminates these occurrences in our experiments.
@@
Our second set of results is shown in Table 3 and in Figure 4. Here the most important impact of utilizing exponentially many linear regions is demonstrated. Not only can more accurate representations of training data be learned, but more linear regions allow the network to better capture underlying nonlinearity to enhance its predictive power in regression tasks. This result is especially significant because it indicates that even in cases where there are fewer data points than linear regions, having the additional regions can still provide performance advantages.
Figure 4: Comparison between standard initialization/gradient descent and pretraining with differentiablity enforced. Using more linear regions allows the curve to better predict the test points.
Table 3: Minimum errors on unseen points from training on sparse data.
Training Type Min $x^{3}$ Min $x^{11}$ Min $\sin(x)$ Min $\tanh(3x)$
Default Network $2.41\times 10^{-4}$ $2.14\times 10^{-3}$ $2.27\times 10^{-5}$ $1.60\times 10^{-4}$
Differentiable Pretraining $\mathbf{5.65\times 10^{-6}}$ $\mathbf{6.53\times 10^{-4}}$ $\mathbf{7.92\times 10^{-7}}$ $\mathbf{5.09\times 10^{-6}}$
@@ 4.3 Gradient Descent does not Directly Optimize Efficiency
Figure 5: Approximation produced with standard methodologies. The layers are shown before ReLU is applied. The neuron colors here are arbitrary; they do not correspond to Figure 2.
Figure 5 shows the interior of a default network. The layers here are shown before applying ReLU. The default networks fail to make efficient use of ReLU to produce linear regions, even falling short of 1 bend per neuron, which can easily be attained by forming a linear spline (1 hidden layer) that interpolates some of the data points. Examining the figure, the first two layers are wasted. No neuron’s activation pattern crosses $y=0$, so ReLU is never used. Layer 3 could be formed directly from the input signal. Deeper in the network, more neurons remain either strictly positive or negative. Those that intersect $y=0$ are monotonic, only able to introduce one bend at a time. The core issue is that while more bends leads to better accuracy, networks that have few bends are not locally connected in parameter space to those that have many. This is problematic since gradients can only carry information about the effects of infinitesimal parameter modifications. If a bend exists, gradient descent can reposition it. But for a neuron that always outputs a strictly positive value (such as the red in layer 2), bends cannot be introduced by infinitesimal weight or bias adjustments. Therefore, bend-related information will be absent from its gradients. Gradient descent will only compel a network to bend by happenstance; indirectly related local factors must guide a neuron to begin outputting negative values. Rarely, these local incentives are totally absent, and the network outputs a bend-free line of best fit.
4.4 Effect of Pretraining on a Compositional Network
@@
In Figure 6 we compare the top performing models with and without reparameterized pretraining. We observe that without the guidance of the pretraining, gradient descent usually loses the triangle generating structure. It typically happens around layer 4 or 5, devolving into noisy patterns and resulting in higher errors, whereas pretraining maintains structure at greater depths. This behavior of gradient descent we observe is problematic since theoretical works often rely on specific constructions within networks to prove their results. Gradient descent greedily abandons any such structure in favor of models that can be worse in the long term. A theoretical result that shows a certain representation exists in the set of neural networks will thus have a hard time actually learning it without a subsequent plan to control training.
@@ 5 Limitations and Extensions
Although our reparemeterization can help maintain the triangular structure out to greater depths than unassisted gradient descent, our neurons will still begin producing non-triangular outputs eventually. This suggests that our reparameterization is limited in the functions it can theoretically capture. It may be enabling a rapid error reduction initially in the first few layers, but be unable to approach most curves to arbitrary precision. It is in this remaining error gap that the second stage, gradient descent, becomes lost. In the appendix, we detail two important results concerning the theoretical expressive power of our reparameterization that support this.
First, the differentiability constraints on the scaling parameters during pretraining necessitate that the network outputs a convex function. It is important for the pretraining to be able to produce a reasonably close initial solution; therefore, the reparameterization we use is only guaranteed to be suitable for approximating convex functions. The second limitation we encounter is that pursuing further smoothness (at least in this 4-neuron configuration) is not possible. The appendix show that approximating $x^{2}$ is the only way to have a twice continuously differentiable output in the infinite depth limit.
Despite these limitations, these networks could still be considered as replacements of the $x^{2}$ gate discussed in Section 2.1. Likewise, they could be used as a base component for assembly into larger networks. These composed networks could learn non-convex functions in higher dimensions, and they would at minimum be guaranteed to represent interpolating polynomials. The drawback is that they would likely become prohibitively large. Rather than approximating another approximation scheme out of ReLU networks, it may be better to focus on finding representations that more naturally arise from ReLU networks themselves when extending this work. Two properties specifically stand out as being somewhat fundamental to deep ReLU approximations: each linear region produced is a full copy of the original input interval $[0,1]$, and the linear regions’ boundaries become dense as depth tends towards infinity. Density ensures that no two inputs are linearly related in an infinite network, and the repeated reproduction of the input space allows each layer to act on each previously existing region independently and simultaneously. These properties both arise in our method, and are likely important considerations when seeking theoretical extensions. Lastly, it might be important to consider non-differentiable model outputs as a way to gain additional flexibility, but additional protections would need to be put in place against overfitting.
@@ 6 Concluding Remarks
This paper focused on exploiting the potential computational complexity advantages neural networks offer for the problem of efficiently learning nonlinear functions; in particular, compelling ReLU networks to approximate functions with exponential accuracy as network depth is linearly increased. Our results showed improvements of one to several orders of magnitude in using our reparameterization and pretraining strategy to train ReLU networks to learn various nonlinear, convex, one-dimensional functions. Though further research is required to extend this technique to general multi-dimensional nonlinear functions, the present work’s principal significance lies in demonstrating the possibility of converting theoretical ReLU-based constructions into novel training procedures. This finding is particularly powerful since random initialization and gradient descent are not likely to produce an efficient solution on their own, even if it can be proven to exist in the set of sufficiently sized ReLU networks. We are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU networks and even more general types of neural networks.